Training Data

Positive Training  (58 cases)
Cottrell precipitator Van de Graaff generator Wimshurst machine aerial  antenna
attenuator  ballast  battery  bimetallic strip  board
brush  capacitance  capacitor  circuit  condenser
control board  control panel  distributer  electric battery  electric cell
electric circuit  electrical circuit  electrical condenser electrical device  electrical distributor
electrical fuse  electrical relay  electrograph  electrostatic generator electrostatic machine
filter  flasher  fuse  inductance  inductor
instrument panel  jack  light ballast  load  plug
precipitator  reactor  rectifier  relay  resistance
security  security measures  security system  solar array  solar battery
solar panel  spark arrester  spark plug  sparking plug  suppresser
transmitting aerial  transponder  zapper
Negative Training  (55 cases)
Andes  Burnett  Diana  DuPonts  Friesland
Gibbs  Hickman  Icarus  Lorraine  Madeira
Quakeress  Southernwood  Waltham  Washington  adventures
affecting  aggrieving  attractiveness  bearer  boll
capitals  concluding  constantly  conviction  damming
deeper  definitions  dimension  discounting  distinctness
exclamation  faking  helplessness  humidly  hurling
introduces  kappa  maims  marine  moderately
monster  parenthesis  pinches  predication  prospect
repudiate  retry  royalty  shopkeepers  soap
sob  swifter  teared  thrashes  tuples
Anchors  (6 dimensions)
bumbled distributor premeditation resistor suppressor

swimmers

Testing Results

Positive tests  Negative tests
Positive  cell, male plug,
Predictions  panel, transducer,

transformer

Negative  Boswellizes, appointer,
Predictions  enforceable, greatness,

planet

Accuracy  10/10 = 100.00%